NPP Projection Admin 2/6
To do this evening/ was completed: A1 2/13 playlists built, old content removed from servers.
Tech'd Peter Hujar, all good. Tech'd Dead Lover too, all good, but we'll need to adjust masking presets to properly cover 1.33 in T3. Sent out a zoom link to the participants of the Q&A plus Lillian and Cristina.
Cable tested in T2 for Super Bowl.
Rescheduled T1 and T4 for today and tomorrow. Cue sheets updated.
Dog Day Afternoon print arrived, in T1.
Presentation Issues:
Preshow Clips/Playlist:
Trailer Updates:
Special Events/ AV set up: Techd and adjusted the playlist for Blazing Saddles, which will be for the private event on Sunday in T6. - LH
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): Ingested the 2D version of Goat, we had the 3D one on the LMS -A
Ingested new content for the week (what was available) for preshow Sundays on Fire V4, Dog Day afternoon preshow, forecasts, etc. Waiting on a few other bits n bobs. - LH
Ingested Kubo and the Two Strings - LH
Outgoing/Shipping (content):
Print/video Inspection: Started inspecting Sundays on Fire print Love Battlefield in T1, which can get finished this evening. Ran reels for Dancing Lady, Coach Carter, Jurassic Park for Cristina/Desmond. - LH
Asset tracking doc updated/checked?:
Maintenance/cleaning: Christie and Dolby had some suggestions for T3 based on the logs we sent, which I carried out. Not sure what effect they may have, if any, but I reported this to DMS so these things can be ruled out if the issue occurs again. -A (Sent the logs/report from this morning - LH)
Outgoing/Shipping (content):
Print/video Inspection: Started inspecting Sundays on Fire print Love Battlefield in T1, which can get finished this evening. Ran reels for Dancing Lady, Coach Carter, Jurassic Park for Cristina/Desmond. - LH
Asset tracking doc updated/checked?:
Maintenance/cleaning: Christie and Dolby had some suggestions for T3 based on the logs we sent, which I carried out. Not sure what effect they may have, if any, but I reported this to DMS so these things can be ruled out if the issue occurs again. -A (Sent the logs/report from this morning - LH)
Also reduced the power to the lamps in T1 and T2 as they were running too high. Only reduced by 15% for flat and 10% for scope. Looked at test patterns and content in both houses and the image is still plenty bright -A
Cue sheets updated/checked (every shift): Yes
Marquee updated?:
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment